OpenStudy (anonymous):

What is the range of the graph of y = 5(x - 2)2 + 7?

OpenStudy (usukidoll):

range is in the y-axis

OpenStudy (anonymous):

so is the range 7<y

OpenStudy (anonymous):

Use graphic calculator to solve: http://www.acalculator.com/graphing-calculator-online/

OpenStudy (anonymous):

The website is kinda confusing, thanks anyway. Can anyone explain how to find the range and what it is for this equation please??

OpenStudy (usukidoll):

you need to graph first then find the range based off that graph which is located in the y-axis

OpenStudy (zzr0ck3r):

The leading coefficient is positive so you know it looks like a smiley face. This is in vertex form with vertex \((-2,7)\). So the lowest value the function takes is \(7\).

OpenStudy (zzr0ck3r):

no need to graph. That is why they gave it in vertex form. A quadratic with positive leading coefficient and vertex \((a,b)\) has range \(y\ge b\) A quadratic with negative leading coefficient and vertex \((a,b)\) has range \(y\le b\)

OpenStudy (anonymous):

Oh ok so y>=7 should be my answer, right?

OpenStudy (usukidoll):

oh... I see. x.x if \[y \geq 7\] then we have a quadratic with a positive leading coefficient and vertex (a,b). IN this case we have a = -2 and b = 7 so (-2,7)

OpenStudy (anonymous):

Yay Thank you everybody who helped